# Poisoning Attacks and Defenses in Recommender Systems: A Survey

### v1

@@ 4.1.2 Targeted Manipulation
Another variant of targeted attacks in poisoning attack is specifically designed to promote certain items to a designated user group, ensuring that these specific users are exposed to the attacker-chosen items. For example, AutoAttack Guo et al. (2023) considers both the target item and user perspectives and it crafts malicious user profiles that closely mimic the attributes and interactions of the targeted user group. This strategy enables an effective influence on the designated group while concurrently minimizing unintended effects on other users. Analogously, UBA Wang et al. (2024) posits that each item has its own unique potential audience and the difficulty of attacking different users varies. A common approach targeting all users can lead to the wasteful expenditure of a fake user budget and a decline in attack efficacy. Therefore, UBA introduces an attack method that considers both performance and budget from a causal perspective, aiming for efficient utilization of resources while maximizing impact to promote targeted items to specific user groups.
4.1.3 Hybrid-Goal Attacks
Attackers might design more flexible attack strategies, aligning with multiple objectives and including both system degradation and targeted manipulation. SGLD Li et al. (2016) stands out as a pioneering method capable of pursuing dual goals concurrently. Its core technique involves maximizing the discrepancy in all predictions before and after the poisoning attack (system degradation) and enhancing the prediction likelihood of a target item for each user (targeted manipulation). It achieves a balance between these two goals using weighted coefficients. Following the groundwork laid by SGLD, studies such as CD-Attack Chen and Li (2019) and NCFAttack Zhang et al. (2020d) have adopted this dual-goal framework. CD-Attack demonstrated that malicious profiles developed using this methodology exhibit good transferability across different domains, while NCFAttack chose to customize a more elegant attack target. It is different from the former, which takes the attack effect as the optimization goal. Instead, it generates specific adversarial gradients based on the model optimization strategy to enable the execution of more precise attacks.
4.2 Attacks Limited by Practical Capability
4.2.1 Knowledge Constraint
In RS model-unknown contexts, attackers often train and employ a surrogate model to deduce representations of RS model, adapting their strategies to the constraints of their knowledge about the victim system. For example, ModelExtractionAttack Yue et al. (2021) employs model distillation techniques, designating the victim black box model as the teacher and the surrogate model as the student. The surrogate model is trained to produce outputs that are as identical as possible to those of the victim black box model. Following the model distillation concept, PC-Attack Zeng et al. (2023) further establishes two objectives: ensuring the representation of the target item is as similar as possible to that of other items, and making the representation of users who interacted with the target item as similar as possible to other users. It then employs a ‘pre-training and fine-tuning’ strategy, where a simulator model is pre-trained on multiple data sources, such as public recommendation datasets. A small portion of the target data is then input into the simulator model for fine-tuning.
In addition to surrogate models, reinforcement learning (RL) Arulkumaran et al. (2017) is also extensively utilized in model-unknown contexts. Attackers design appropriate reward functions for RL to optimize the attack effect without directly using the victim model to provide simulated representations of RS. For example, LOKI Zhang et al. (2020b) implements RL through designing the reward as the weighted averaged influence on the prediction scoring function, facilitating the generation of user behavior samples, while CopyAttack Fan et al. (2021) applies RL based on a ranking scoring function that will assign a positive reward if the target item belongs to the recommended list, to identify and replicate real user profiles from a source domain into the target domain.
4.2.2 Cost Limitation
4.2.3 Invisibility
4.3 Attacks Specific to Victim Architecture
4.3.1 Centralized Recommendation
Starting with traditional decentralized systems, studies such as TNAFang et al. (2020) explore the vulnerabilities of matrix factorization-based RS, while UNAttackChen et al. (2021) investigates neighborhood-based systems. Both methods construct hypothetical rating matrices post-malicious file injection and iteratively optimize these matrices to evaluate and identify more threatening attack vectors. As research progresses, it extends to more complex deep learning-based RS. DLAttackHuang et al. (2021) targets pure deep learning-based RS, and NCFAttackZhang et al. (2020d) focuses on neural collaborative filtering-based RS. Both approaches develop rigorous models to specify attack constraints within their recommendation architecture and define various objective functions to capture the fundamental goals of availability and targeted attacks. Furthermore, GOAT Wu et al. (2021d) and GSPAttack Nguyen et al. (2022) execute poisoning attacks specifically on graph-based RS that predominantly incorporate GNN Zhang et al. (2019a) in their architectures. These studies expose vulnerabilities in the gradient update mechanisms of these DL-based systems. This evolution continues with the advent of Self-Supervised Learning (SSL)  Yu et al. (2023a). SSLAttack Wang et al. (2023a) targets this paradigm, shifting the focus of the poisoning attacks to the initial pre-training phase. This approach stresses the injection of adversarial influences during pre-training, which are subsequently propagated to downstream recommendation models during the fine-tuning stage. In a similar context, PromptAttack Wu et al. (2023c) explores vulnerabilities in pre-trained models utilizing prompt learning Zhou et al. (2022), a nuanced approach compared to traditional fine-tuning, thus exposing weaknesses in various downstream recommendation models. CLeaR Wang et al. (2023d) also emphasizes the focus on SSL, intensifying contrastive learning loss in RS can result in a uniform distribution of representations. Additionally, in response to the growing demand for transparency, explainable RS, which converts user interactions into logical expressions, has been scrutinized. TextRecAttack Zhang et al. (2024a) highlights the security risks associated with integrating large language models (LLMs) Fan et al. (2023b) into RS. This attack illustrates how minor textual modifications during the test phase can significantly enhance an item’s exposure, circumventing the need for direct manipulation of the model’s training infrastructure. H-CARS Chen et al. (2023b) leverages counterfactual explanations to craft malicious user profiles, exploiting the logical reasoning in these systems and thus uncovering new vulnerabilities.
4.3.2 Decentralized Recommendation
As concerns over privacy in centralized systems grew, research pivoted towards decentralized federated RS, where local storage of user data enhances security and privacy. Despite the enhanced robustness of this architecture, FedRecAttack Rong et al. (2022b) illustrates that public interactions in federated RS can be exploited to approximate user features. These derived user features can subsequently be utilized to intensify the manipulation of recommendations. PSMU Yuan et al. (2023a) identifies a high similarity between the top-K recommendations for randomly generated synthetic users and actual users. This observation indicates that attackers can manipulate synthetic users to obtain feedback on targeted items, and this feedback can subsequently be leveraged to refine the attack model. Building on the above work, A-hum Rong et al. (2022a) further focuses on identifying ‘hard users’ who consider the target items as negative samples and manipulate malicious interactions to elevate these users’ preferences for the targeted items. PipAttack Zhang et al. (2022a) capitalizes on the distinctive characteristics of popular item aggregation within the widely used BPR loss. It alters the target item to resemble the attributes of popular items in the embedding space, thereby increasing the likelihood of user interactions with that item. UA-FedRec Yi et al. (2023) modifies the representations to increase the distance between similar samples and decrease it for dissimilar ones to standardize the distances across all user and item representations. This deliberately created uniform distribution can disrupt each user’s recommendation results and degrade overall recommendation performance.
4.4 Attacks Divided by Poisoning Strategies
4.4.1 Data Poisoning
In scenarios involving multimodal data, the specific vulnerabilities to poisoning attacks become more evident. PSMU(V) Yuan et al. (2023c) and IPDGI Chen et al. (2023a) use guided diffusion models to create adversarial images aimed at facilitating item promotion. These models capably emulate the distribution of benign images, resulting in adversarial images that closely resemble the original ones in terms of fidelity. Furthermore, methods like TDP-CP Zhang et al. (2022b) and ARG Chiang et al. (2023) concentrate on the manipulation of textual data associated with items. They implement methods like adding, deleting, or replacing words or even generating new textual content with the objective of inducing shifts in recommendations. More modal information is incorporated with the introduction of KGs Ji et al. (2021). These elements add detailed entity attributes, such as user demographics and item descriptions. KGAttack Chen et al. (2022), and KG-RLattack Wu et al. (2022b) utilize the extensive information in KGs to create synthetic user profiles that are both authentic and credible. These poisoning attack methods illustrate the extensive range of poisoning attacks across different data modalities.
4.4.2 Model Poisoning
Up to now, model poisoning attacks have mainly appeared in decentralized RS. For example, PipAttack Zhang et al. (2022a) uploads carefully crafted gradients via a small number of malicious users during the model update to enable the target item to have the characteristics of popular items in the embedding space. FedRecAttack Rong et al. (2022b) makes use of the public interactions to approximate users’ feature vectors, then design a bi-level optimization technique to generate poisoned gradients accordingly and control malicious users to upload the poisoned gradients in a well-designed way. A-hum Rong et al. (2022a) optimizes the presentation by gradient descent to mine hard users for a target item, then upload poisoned gradient into the server. FedAttack Wu et al. (2022a) adopts a strategy in which it selects items that are most and least relevant to user embeddings as the hardest negative and positive samples, respectively. Malicious clients then generate and upload poisoned gradients designed to swap the model’s predictions for these two samples, thereby confusing the model’s output.
Remark. Data poisoning attacks and model poisoning attacks exhibit clear differences in their approaches and their effects on the system, making it relatively straightforward to choose between them based on the attack objective. However, it is important to note that model poisoning attacks require the uploading of poisoned gradients, making them primarily infeasible in centralized recommender systems. On the other hand, data poisoning attacks can occur in both decentralized and centralized scenarios. For instance, PSMU(V)  Yuan et al. (2023c) involves constructing and uploading adversarial images poisoned by synthetic malicious users in a centralized recommendation setting.
5 Defense against Poisoning Attacks
5.1 Poisoning Data Filtering
5.1.1 Supervised Anomaly Filtering
5.1.2 Unsupervised Clustering Filtering
Unsupervised clustering filtering methods are particularly valuable in scenarios where labels are scarce, employing techniques that can discern patterns and group similar data points without prior categorization. One mainstream is to combine malicious context hidden in recommendations into improved K-Means algorithms. For example, ARBD Yu et al. (2019) utilizes a modified K-Means algorithm that leverages two specific metrics: the popularity degree of an item and the distance between users. By adjusting the clustering criteria to these particular dimensions, ARBD enhances its ability to detect outlier activities more effectively. GAGE Zhang et al. (2020a) extends the utility of clustering by constructing a user relationship graph based on user rating behaviors. It employs graph embedding techniques to transform the complex relationships into low-dimensional vector representations of each node (user). Following this transformation, the K-Means++ clustering algorithm is applied to these vectors to form groups and facilitate the identification of malicious activities. GSD Zhang and Wang (2020) introduces a temporal element into the clustering process by incorporating time sequential information. GSD analyzes the scoring trajectory of each item, segmenting these trajectories based on fixed time intervals to form candidate groups. Then the bisection k-Means algorithm is used to further classify these groups based on their degree of suspicion, effectively isolating the groups likely involved in shilling attacks.
Aside from the well-known K-Means algorithm, several other clustering techniques have demonstrated their effectiveness in detecting anomalies within various systems. These methods leverage diverse approaches to clustering, utilizing spectral clustering, density-based clustering, and even hierarchical clustering to enhance detection capabilities. BS-SC Cai and Zhang (2019a) explores the capabilities of spectral clustering. This method manually defines multiple types of features to construct a user behavior similarity matrix and then reconstructs the behavior graph based on rank computation. By utilizing the spectral clustering algorithm, it effectively clusters shilling profiles based on their behavioral similarities. DSA-AURB Cai and Zhang (2019b) employs a density-based clustering approach. Initially, it identifies target items and the intentions of attacking users by analyzing deviations in rating patterns for each item. The analysis extends to users’ rating behaviors from an interest preference perspective, using entropy and block entropy metrics to measure the diversity and memory of users’ preferences. These combined metrics are then used as the density information for clustering.
Another approach involves enhancing basic clustering algorithms with additional components to improve clustering performance. For instance, MMD Xu et al. (2020) integrates metric learning and clustering. It begins by converting reviews into sentiment scores, followed by a sophisticated malicious user profiling component that identifies discrepancies between sentiment scores and ratings. This method incorporates these discrepancies to derive a suitable metric matrix, which is then added to the clustering approach for detection. Similarly, UD-HMM Zhang et al. (2018) combines a hidden Markov model (HMM) with hierarchical clustering to detect shilling attacks. It models users’ historical rating behaviors using HMM and evaluates each user’s level of suspicion by analyzing their preference sequences and rating behavior differences compared to genuine users. A hierarchical clustering method then groups users based on their suspicious degrees.
5.1.3 Confidence-Based Filtering
Building on these approaches, some methods further explore unique scenarios to address specific challenges or new conditions. For instance, PDR Lai et al. (2023) aims to address the problem of noisy data labels, encompassing both benign noise and malicious poisoned data. This method employs a strategy that dynamically adjusts prior knowledge to refine and enhance the detection of malicious profiles. This adaptive approach allows PDR to continuously update and improve its detection capabilities as more data becomes available or as user behaviors evolve. Additionally, FSAD Jiang et al. (2020) pioneers the study of shilling attacks within the context of federated learning. FSAD innovatively designs four novel features derived from the gradients exchanged among clients during the learning process. By leveraging these gradient-based features, FSAD trains a semi-supervised Bayes classifier using an EM algorithm to effectively identify shilling attackers.
5.2 Robust Training
Model Main Technologies Strategy
TSA-TF Xu and Zhang (2019) SVM Supervised Anomaly Filtering
MMD Xu et al. (2020) Metric Learning Unsupervised Clustering Filtering
PRD-BGP Liu (2020) Pairwise Learning Confidence-Based Filtering
5.2.1 Collaborative Enhancement Training
5.2.2 Confidence-Aware Training
The concept of reweighted loss, aimed at adaptively managing the influence of training data, has been developed and diversified through several innovative methods. DeCA Zhang et al. (2023) observes that different models tend to make similar predictions on clean examples, which reflect real user preferences, whereas predictions on noisy examples vary more significantly. Motivated by this observation, DeCA minimizes the KL divergence between the user preference distributions parameterized by two recommendation models. It simultaneously maximizes the likelihood of data observation, thereby dynamically adjusting weights based on each data point’s assessed noise level. Analogously, SGDL Gao et al. (2022) collected memorized interactions identified during the early training stage and used them as denoising signals. These interactions guide the later ‘noise-sensitive’ phases of the model’s training in a meta-learning fashion. BOD Wang et al. (2023c) models recommendation denoising as a bi-level optimization problem. It leverages the denoising gradient information generated from the inner optimization process to guide weight determination in the outer optimization for the training data. AutoDenoise Lin et al. (2023) addresses the challenge of highly dynamic data distributions by employing a deep RL-based framework. It incorporates an instance-denoising policy network that adaptively selects noise-free and predictive data instances.
5.2.3 Contextual Enrichment Training
6 Evaluation
6.1 Attack Metric
[…]
    $\text{Precision}_{rec}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}{\frac{|\mathcal{I}_{u}\cap\mathcal{I}^{rec}_{u}|}{\mathcal{I}^{rec}_{u}}},$ (13)
    where $\mathcal{I}_{u}$ represents the actual items interacted by user $u$.
    $\text{Recall}_{rec}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}{\frac{|\mathcal{I}_{u}\cap\mathcal{I}^{rec}_{u}|}{\mathcal{I}_{u}}},$ (14)
6.2 Defense Metric
[…]
Robust Training Metric. This group concentrates on assessing the overall performance of the recommendation system. Consequently, the key metrics employed are akin to those used in evaluating untargeted attack metrics, including MAE, RMSE, $\text{Precision}_{rec}$, and $\text{Recall}_{rec}$, referred to Equation 11121314.

### v2

@@ 4.1.2 Targeted Manipulation
Another variant of of targeted attacks in poisoning involves the deliberate promotion of certain items to a specific group of users. This strategy ensures that these chosen users encounter the items selected by the attacker. For example, AutoAttack Guo et al. (2023) considers both the target item and user perspectives and it crafts malicious user profiles that closely mimic the attributes and interactions of the targeted user group. This strategy enables an effective influence on the designated group while concurrently minimizing unintended effects on other users. Analogously, UBA Wang et al. (2024) proposes that each item has a distinct target audience and that the challenge of attacking various users differs. A common approach targeting all users can lead to the wasteful expenditure of a fake user budget and a decline in attack efficacy. Therefore, UBA introduces an attack method that considers both performance and budget from a causal perspective, aiming for efficient utilization of resources while maximizing impact to promote targeted items to specific user groups.
@@
4.1.3 Hybrid-Goal Attacks
Attackers might design more flexible attack strategies, aligning with multiple objectives and including both system degradation and targeted manipulation. SGLD Li et al. (2016) stands out as a pioneering method capable of pursuing dual goals concurrently. Its core technique involves maximizing the discrepancy in all predictions before and after the poisoning attack (system degradation) and enhancing the prediction likelihood of a target item for each user (targeted manipulation). It achieves a balance between these two goals using weighted coefficients. Following the groundwork laid by SGLD, studies such as CD-Attack Chen and Li (2019) and NCFAttack Zhang et al. (2020d) have adopted this dual-goal framework. CD-Attack demonstrates that malicious profiles developed using this methodology exhibit good transferability across different domains, while NCFAttack chooses to customize a more elegant attack target. It is different from the former, which takes the attack effect as the optimization goal. Instead, it generates specific adversarial gradients based on the model optimization strategy to enable the execution of more precise attacks.
4.2 Attacks Limited by Practical Capability
4.2.1 Knowledge Constraint
In RS model-unknown contexts, attackers often train and employ a surrogate model to deduce representations of RS model, adapting their strategies to the constraints of their knowledge about the victim system. For example, ModelExtractionAttack Yue et al. (2021) employs model distillation techniques, designating the victim black box model as the teacher and the surrogate model as the student. The surrogate model is trained to produce outputs that are as identical as possible to those of the victim black box model. Following the model distillation concept, PC-Attack Zeng et al. (2023) further establishes two objectives: ensuring the representation of the target item is as similar as possible to that of other items, and making the representation of users who interacted with the target item as similar as possible to other users. It then employs a “pre-training and fine-tuning” strategy, where a simulator model is pre-trained on multiple data sources, such as public recommendation datasets. A small portion of the target data is then input into the simulator model for fine-tuning.
In addition to surrogate models, reinforcement learning (RL) Arulkumaran et al. (2017) is also extensively utilized in model-unknown contexts. Attackers design appropriate reward functions for RL to optimize the attack effect without directly using the victim model to provide simulated representations of RS. For example, LOKI Zhang et al. (2020b) implements RL through designing the reward as the weighted averaged influence on the prediction scoring function, facilitating the generation of user behavior samples, while CopyAttack Fan et al. (2021) employs RL using a ranking scoring function, which grants a positive reward if the target item appears in the recommended list. This method identifies and replicates real user profiles from a source domain into the target domain.
4.2.2 Cost Limitation
4.2.3 Invisibility
4.3 Attacks Specific to Victim Architecture
4.3.1 Centralized Recommendation
Starting with traditional centralized systems, studies such as TNAFang et al. (2020) explore the vulnerabilities of matrix factorization-based RS, while UNAttackChen et al. (2021) investigates neighborhood-based systems. Both methods construct hypothetical rating matrices post-malicious file injection and iteratively optimize these matrices to evaluate and identify more threatening attack vectors. As research progresses, it extends to more complex deep learning-based RS. DLAttackHuang et al. (2021) targets pure deep learning-based RS, and NCFAttackZhang et al. (2020d) focuses on neural collaborative filtering-based RS. Both approaches develop rigorous models to specify attack constraints within their recommendation architecture and define various objective functions to capture the fundamental goals of availability and targeted attacks. Furthermore, GOAT Wu et al. (2021d) and GSPAttack Nguyen et al. (2022) execute poisoning attacks specifically on graph-based RS that predominantly incorporate GNN Zhang et al. (2019a) in their architectures. These studies expose vulnerabilities in the gradient update mechanisms of these DL-based systems. This evolution continues with the advent of Self-Supervised Learning (SSL)  Yu et al. (2023a). SSLAttack Wang et al. (2023a) targets this paradigm, shifting the focus of the poisoning attacks to the initial pre-training phase. This approach stresses the injection of adversarial influences during pre-training, which are subsequently propagated to downstream recommendation models during the fine-tuning stage. In a similar context, PromptAttack Wu et al. (2023c) explores vulnerabilities in pre-trained models utilizing prompt learning Zhou et al. (2022), a nuanced approach compared to traditional fine-tuning, thus exposing weaknesses in various downstream recommendation models. CLeaR Wang et al. (2023d) also emphasizes the focus on SSL, intensifying contrastive learning loss in RS can result in a uniform distribution of representations. Additionally, in response to the growing demand for transparency, explainable RS, which converts user interactions into logical expressions, has been scrutinized. TextRecAttack Zhang et al. (2024a) highlights the security risks associated with integrating large language models (LLMs) Fan et al. (2023b) into RS. This attack illustrates how minor textual modifications during the test phase can significantly enhance an item’s exposure, circumventing the need for direct manipulation of the model’s training infrastructure. H-CARS Chen et al. (2023b) leverages counterfactual explanations to craft malicious user profiles, exploiting the logical reasoning in these systems and thus uncovering new vulnerabilities.
4.3.2 Decentralized Recommendation
As concerns over privacy in centralized systems grew, research pivoted towards decentralized federated RS, where local storage of user data enhances security and privacy. Despite the enhanced robustness of this architecture, FedRecAttack Rong et al. (2022b) illustrates that public interactions in federated RS can be exploited to approximate user features. These derived user features can subsequently be utilized to intensify the manipulation of recommendations. PSMU Yuan et al. (2023a) identifies a high similarity between the top-K recommendations for randomly generated synthetic users and actual users. This observation indicates that attackers can manipulate synthetic users to obtain feedback on targeted items, and this feedback can subsequently be leveraged to refine the attack model. Building on the above work, A-hum Rong et al. (2022a) further focuses on identifying hard users who consider the target items as negative samples and manipulate malicious interactions to elevate these users’ preferences for the targeted items. PipAttack Zhang et al. (2022a) capitalizes on the distinctive characteristics of popular item aggregation within the widely used BPR loss. It alters the target item to resemble the attributes of popular items in the embedding space, thereby increasing the likelihood of user interactions with that item. UA-FedRec Yi et al. (2023) modifies the representations to increase the distance between similar samples and decrease it for dissimilar ones to standardize the distances across all user and item representations. This deliberately created uniform distribution can disrupt each user’s recommendation results and degrade overall recommendation performance.
4.4 Attacks Divided by Poisoning Strategies
4.4.1 Data Poisoning
In scenarios involving multimodal data, the specific vulnerabilities to poisoning attacks become more evident. PSMU(V) Yuan et al. (2023c) and IPDGI Chen et al. (2023a) use guided diffusion models to create adversarial images aimed at facilitating item promotion. These models capably emulate the distribution of benign images, resulting in adversarial images that closely resemble the original ones in terms of fidelity. Furthermore, methods like TDP-CP Zhang et al. (2022b) and ARG Chiang et al. (2023) concentrate on the manipulation of textual data associated with items. They implement methods like adding, deleting, replacing words, or even generating new textual content with the objective of inducing shifts in recommendations. More modal information is incorporated with the introduction of KGs Ji et al. (2021). These elements add detailed entity attributes, such as user demographics and item descriptions. KGAttack Chen et al. (2022) and KG-RLattack Wu et al. (2022b) utilize the extensive information in KGs to create synthetic user profiles that are both authentic and credible. These poisoning attack methods illustrate the extensive range of poisoning attacks across different data modalities.
4.4.2 Model Poisoning
Up to now, model poisoning attacks have mainly appeared in decentralized RS. For example, PipAttack Zhang et al. (2022a) involves a few malicious users uploading meticulously crafted gradients during the model’s update process. This allows the target item to adopt the attributes of popular items within the embedding space. FedRecAttack Rong et al. (2022b) makes use of the public interactions to approximate users’ feature vectors, and then design a bi-level optimization technique to generate poisoned gradients accordingly and control malicious users to upload the poisoned gradients in a well-designed way. A-hum Rong et al. (2022a) optimizes the presentation by gradient descent to mine hard users for a target item, and then uploads poisoned gradient into the server. FedAttack Wu et al. (2022a) adopts a strategy in which it selects items that are most and least relevant to user embeddings as the hardest negative and positive samples, respectively. Malicious clients then generate and upload poisoned gradients designed to swap the model’s predictions for these two samples, thereby confusing the model’s output.
Remark. Data poisoning attacks and model poisoning attacks exhibit clear differences in their approaches and their effects on the system, making it relatively straightforward to choose between them based on the attack objective. However, it is important to note that model poisoning attacks require the uploading of poisoned gradients, making them primarily infeasible in centralized recommender systems. On the other hand, data poisoning attacks can occur in both decentralized and centralized scenarios. For instance, PSMU(V)  Yuan et al. (2023c) involves constructing and uploading adversarial images poisoned by synthetic malicious users in a decentralized recommendation setting.
5 Defense against Poisoning Attacks
5.1 Poisoning Data Filtering
5.1.1 Supervised Anomaly Filtering
5.1.2 Unsupervised Clustering Filtering
Unsupervised clustering filtering methods are particularly valuable in scenarios where labels are scarce, employing techniques that can discern patterns and group similar data points without prior categorization. One mainstream is to combine malicious context hidden in recommendations into improved K-Means algorithms. For example, ARBD Yu et al. (2019) utilizes a modified K-Means algorithm that leverages two specific metrics: the popularity degree of an item and the distance between users. By adjusting the clustering criteria to these particular dimensions, ARBD enhances its ability to detect outlier activities more effectively. GAGE Zhang et al. (2020a) extends the utility of clustering by constructing a user relationship graph based on user rating behaviors. It employs graph embedding techniques to transform the complex relationships into low-dimensional vector representations of each node (user). Following this transformation, the K-Means++ clustering algorithm is applied to these vectors to form groups and facilitate the identification of malicious activities. GSD Zhang and Wang (2020) introduces a temporal element into the clustering process by incorporating time sequential information. GSD analyzes the scoring trajectory of each item, segmenting these trajectories based on fixed time intervals to form candidate groups. Then the bisection k-Means algorithm is used to further classify these groups based on their degree of suspicion, effectively isolating the groups likely involved in poisoning attacks.
Aside from the well-known K-Means algorithm, several other clustering techniques have demonstrated their effectiveness in detecting anomalies within various systems. These methods leverage diverse approaches to clustering, utilizing spectral clustering, density-based clustering, and even hierarchical clustering to enhance detection capabilities. BS-SC Cai and Zhang (2019a) explores the capabilities of spectral clustering. This method manually defines multiple types of features to construct a user behavior similarity matrix and then reconstructs the behavior graph based on rank computation. By utilizing the spectral clustering algorithm, it effectively clusters malicious profiles based on their behavioral similarities. DSA-AURB Cai and Zhang (2019b) employs a density-based clustering approach. Initially, it identifies target items and the intentions of attacking users by analyzing deviations in rating patterns for each item. The analysis extends to users’ rating behaviors from an interest preference perspective, using entropy and block entropy metrics to measure the diversity and memory of users’ preferences. These combined metrics are then used as the density information for clustering.
Another approach involves enhancing basic clustering algorithms with additional components to improve clustering performance. For instance, MMD Xu et al. (2020) integrates metric learning and clustering. It begins by converting reviews into sentiment scores, followed by a sophisticated malicious user profiling component that identifies discrepancies between sentiment scores and ratings. This method incorporates these discrepancies to derive a suitable metric matrix, which is then added to the clustering approach for detection. Similarly, UD-HMM Zhang et al. (2018) combines a hidden Markov model (HMM) with hierarchical clustering to detect poisoning attacks. It models users’ historical rating behaviors using HMM and evaluates each user’s level of suspicion by analyzing their preference sequences and rating behavior differences compared to genuine users. A hierarchical clustering method then groups users based on their suspicious degrees.
5.1.3 Confidence-Based Filtering
Building on these approaches, some methods further explore unique scenarios to address specific challenges or new conditions. For instance, PDR Lai et al. (2023) aims to address the problem of noisy data labels, encompassing both benign noise and malicious poisoned data. This method employs a strategy that dynamically adjusts prior knowledge to refine and enhance the detection of malicious profiles. This adaptive approach allows PDR to continuously update and improve its detection capabilities as more data becomes available or as user behaviors evolve. Additionally, FSAD Jiang et al. (2020) pioneers the study of malicious attacks within the context of federated learning. FSAD innovatively designs four novel features derived from the gradients exchanged among clients during the learning process. By leveraging these gradient-based features, FSAD trains a semi-supervised Bayes classifier using an EM algorithm to effectively identify malicious attackers.
5.2 Robust Training
Model Main Technologies Strategy
TSA-TF Xu and Zhang (2019) SVM Supervised Anomaly Filtering
MMD Xu et al. (2020) Metric Learning Unsupervised Clustering Filtering
PRD-BGP Liu (2020) Pairwise Learning Confidence-Based Filtering
5.2.1 Collaborative Enhancement Training
5.2.2 Confidence-Aware Training
The concept of reweighted loss, aimed at adaptively managing the influence of training data, has been developed and diversified through several innovative methods. DeCA Zhang et al. (2023) notes that various models generally produce similar predictions when dealing with clean examples, which reflect real user preferences, whereas predictions on noisy examples vary more significantly. Motivated by this observation, DeCA minimizes the KL divergence between the user preference distributions parameterized by two recommendation models. It simultaneously maximizes the likelihood of data observation, thereby dynamically adjusting weights based on each data point’s assessed noise level. Analogously, SGDL Gao et al. (2022) collects memorized interactions identified during the early training stage and uses them as denoising signals. These interactions guide the later “noise-sensitive” phases of the model’s training in a meta-learning fashion. BOD Wang et al. (2023c) models recommendation denoising as a bi-level optimization problem. It leverages the denoising gradient information generated from the inner optimization process to guide weight determination in the outer optimization for the training data. AutoDenoise Lin et al. (2023) addresses the challenge of highly dynamic data distributions by employing a deep RL-based framework. It incorporates an instance-denoising policy network that adaptively selects noise-free and predictive data instances.
5.2.3 Contextual Enrichment Training
6 Evaluation
6.1 Attack Metric
[…]
    $\text{Precision}_{rec}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}{\frac{|\mathcal{I}_{u}\cap\mathcal{I}^{rec}_{u}|}{\mathcal{I}^{rec}_{u}}}.$ (13)
    where $\mathcal{I}_{u}$ represents the actual items interacted by user $u$.
    $\text{Recall}_{rec}=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}{\frac{|\mathcal{I}_{u}\cap\mathcal{I}^{rec}_{u}|}{\mathcal{I}_{u}}}.$ (14)
6.2 Defense Metric
[…]
Robust Training Metric. This group concentrates on assessing the overall performance of the recommendation system. Consequently, the key metrics employed are akin to those used in evaluating untargeted attack metrics, including MAE, RMSE, $\text{Precision}_{rec}$, and $\text{Recall}_{rec}$, referred to Equations 111213, and 14.